# Scalable Diffusion Models with State Space Backbone

### v2

@@ 3.2 Model Analysis
We investigate scaling properties of DiS by studying the effect of depth, i.e., number of SSM layers, width, e.g. the hidden size. Specifically, we train 5 DiT models on the ImageNet dataset with a resolution of 256$\times$256, spanning model configurations from small to huge as detailed in Table 1, denoted as (S, B, M, L, H) for simple. As shown in Figure 3 (b), the performance improves as the depth increase from 25 to 49. Similarly, increasing the width from 387 to 768 yields performance gains. Notably, the model’s performance has not yet reached convergence. Overall, across all five configurations, significant enhancements in FID are observed across all training stages by augmenting the depth and width of the SSM structure.
@@ 3.4 Case Study
Table 5: Benchmarking class-conditional image generation on ImageNet 512$\times$512. DiT-H/2 demonstrates a promising performance compared with both CNN-based and Transformer-based UNet for diffusion.
@@ 4.1 State Space Backbone for Sequence Modeling
State space models, a recent addition to the realm of deep learning, have garnered attention for their capacity to transform state spaces [20, 19, 31]. Drawing inspiration from continuous state space models in control systems and leveraging the HiPPO initialization [17], LSSL [20] has demonstrated promise in addressing long-range dependency issues. However, the computational and memory demands associated with the state representation render LSSL impractical for real-world applications. To mitigate this challenge, S4 [19] proposes normalizing parameters into a diagonal structure. Subsequently, various iterations of structured state-space models have emerged, featuring diverse structures such as complex-diagonal configurations [22, 18], multiple-input multiple-output supporting [59], decomposition of diagonal plus low-rank operations [23], selection mechanism [16]. These models are then integrated into large representation models [44, 43, 15]. These models have been integrated into large representation models [44, 43, 15], primarily focusing on applications involving long-range and causal data such as language and speech, including tasks like language understanding [43] and content-based reasoning [16]. In this paper, we study the scaling properties of SSM when used as the backbone of diffusion models of images.
@@ 4.2 Image Generation in Diffusions
Diffusion models [60, 7, 9, 26, 48] represent a category of generative probabilistic models designed to approximate data distributions and facilitate straightforward sampling procedures. Typically, these models operate by taking a Gaussian noise input and iteratively denoising it through a series of gradual steps until it conforms to the target distribution. The specifics of the denoising process, including the number of steps and the transformation parameterization, vary across different studies [60, 61, 42, 41, 39]. Recently, diffusion models have emerged as leading generators renowned for their ability to learn intricate distributions and produce diverse, high-quality samples. These models have found successful application across various domains, including images [9, 48, 14, 55, 53, 54, 13, 12], video [58], 3D scenes [46], and motion sequences [68, 65]. In this work, we investigate novel structural enhancements for diffusion implementation. Here we explore a new structure class for diffusion implementation.

### v3

@@ Abstract
Keywords:
State space diffusion models image synthesis
@@ 3.2 Model Analysis
We investigate scaling properties of DiS by studying the effect of depth, i.e., number of SSM layers, width, e.g. the hidden size. Specifically, we train 5 DiT models on the ImageNet dataset with a resolution of 256$\times$256, spanning model configurations from small to huge as detailed in Table 1, denoted as (S, B, M, L, H) for simple. As shown in Figure 3 (b), the performance improves as the depth increase from 25 to 49. Similarly, increasing the width from 384 to 768 yields performance gains. Notably, the model’s performance has not yet reached convergence. Overall, across all five configurations, significant enhancements in FID are observed across all training stages by augmenting the depth and width of the SSM structure.
@@ 3.4 Case Study
Table 5: Benchmarking class-conditional image generation on ImageNet 512$\times$512. DiS-H/2 demonstrates a promising performance compared with both CNN-based and Transformer-based UNet for diffusion.
@@ 4.1 State Space Backbone for Sequence Modeling
State space models, a recent addition to the realm of deep learning, have garnered attention for their capacity to transform state spaces [20, 19, 31]. Drawing inspiration from continuous state space models in control systems and leveraging the HiPPO initialization [17], LSSL [20] has demonstrated promise in addressing long-range dependency issues. However, the computational and memory demands associated with the state representation render LSSL impractical for real-world applications. To mitigate this challenge, S4 [19] proposes normalizing parameters into a diagonal structure. Subsequently, various iterations of structured state-space models have emerged, featuring diverse structures such as complex-diagonal configurations [22, 18], multiple-input multiple-output supporting [59], decomposition of diagonal plus low-rank operations [23], selection mechanism [16]. These models have been integrated into large representation models [44, 43, 15], primarily focusing on applications involving long-range and causal data such as language and speech, including tasks like language understanding [43] and content-based reasoning [16]. Seminal work [67] is most related to us, they analyze attention-free structure for image synthesis. In this paper, we focus on the scaling properties of advanced Mamba structure when used as the backbone of diffusion models of images.
@@ 4.2 Image Generation in Diffusions
Diffusion models [60, 7, 9, 26, 48] represent a category of generative probabilistic models designed to approximate data distributions and facilitate straightforward sampling procedures. Typically, these models operate by taking a Gaussian noise input and iteratively denoising it through a series of gradual steps until it conforms to the target distribution. The specifics of the denoising process, including the number of steps and the transformation parameterization, vary across different studies [60, 61, 42, 41, 39]. Recently, diffusion models have emerged as leading generators renowned for their ability to learn intricate distributions and produce diverse, high-quality samples. These models have found successful application across various domains, including images [9, 48, 14, 55, 53, 54, 13, 12], video [58], 3D scenes [46], motion sequences [69, 65], and so on. We follow the path of image generation with diffusion, but consider the impact of high-efficient structures.